# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v3

@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\sigma^{2}_{Y_{k}}\end{bmatrix}-\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\boldsymbol{{\Sigma}_{Z+\epsilon}^{-1}}\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\mathrm{cov}^{2}(Y_{1},Z_{1}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{1}}}-\frac{\lambda}{\sigma^{4}_{Z_{1}}})&...&\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ \vdots&&\vdots\\ \mathrm{cov}(Y_{k},Z_{k}+\epsilon)\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&...&\sigma^{2}_{Y_{k}}-\mathrm{cov}^{2}(Y_{k},Z_{k}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{k}}}-\frac{\lambda}{\sigma^{4}_{Z_{k}}})\\ \end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&&\\ &\ddots&\\ &&\sigma^{2}_{Y_{k}}-\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}+\lambda\begin{bmatrix}\frac{1}{\sigma^{4}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{1},Z_{1})\mathrm{cov}(Y_{k},Z_{k})\\ \vdots&&\vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{k},Z_{k})\mathrm{cov}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{4}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}\right|\end{split}$
@@
We use the notation $\boldsymbol{v}=\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}^{T}$, and $\boldsymbol{V}=\mathrm{diag}(\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1}),\cdots,\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k}))$. And utilize the rule of determinants of sums [36], then we have:
$\begin{split}|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|=&|(\boldsymbol{\Sigma_{Y}-V})+\lambda\boldsymbol{v}\boldsymbol{v}^{T}|\\ =&|\boldsymbol{\Sigma_{Y}-V}|+\lambda\boldsymbol{v}^{T}\boldsymbol{(\boldsymbol{\Sigma_{Y}-V})^{*}}\boldsymbol{v}\end{split}$ (53)
where $(\boldsymbol{\Sigma_{Y}-V})^{*}$ is the adjoint of the matrix $(\boldsymbol{\Sigma_{Y}-V})$. For simplicity, we can rewrite ${|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$ as:
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})+\Omega\end{split}$ (54)
@@
where $\Omega=\lambda\boldsymbol{v}^{T}\boldsymbol{(\boldsymbol{\Sigma_{Y}-V})}^{*}\boldsymbol{v}$. The specific value of $\Omega$ can be obtained as:
@@
$\begin{split}\Omega=\lambda\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\begin{bmatrix}V_{11}&&\\ &\ddots&\\ &&V_{kk}\end{bmatrix}\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})\\ \vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\end{split}$ (55)
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})\end{split}$ (58)
Accordingly, $|\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|$ is:
$|\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|=\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))$ (59)
As a result, $\frac{|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}{|\boldsymbol{\Sigma_{Y|Z}}|}$ is expressed as:
$\frac{|\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}=\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})}$ (60)
Combine Equations 60 and 50 together, the entropy change is expressed as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (61)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&\mathrm{if}\ M>0\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\mathrm{if}\ M\leq 0\end{matrix}\right.$ (64)
@@ B.2 Add Gaussian Noise to Raw Images
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X+\epsilon})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X+\epsilon}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\end{split}$ (66)
@@
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}$ (67)
Clearly, the introduction of Gaussian noise to each pixel in the original images has a detrimental impact on the task. Note that some studies have empirically shown that adding Gaussian noise to partial pixels of input images may be beneficial to the learning task [30] [69].
@@ C.1 Inject Linear Transform Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})=&H(\mathcal{T};\boldsymbol{Z})-H(\mathcal{T};\boldsymbol{Z}+Q\boldsymbol{Z})\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y};\boldsymbol{Z}+Q\boldsymbol{Z})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Y};\boldsymbol{Z}+Q\boldsymbol{Z})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma}_{\boldsymbol{Z}}||\boldsymbol{\Sigma_{Y}}-\boldsymbol{\Sigma_{YZ}}{\boldsymbol{\Sigma_{Z}^{-1}}}\boldsymbol{\Sigma_{ZY}}|}{|\boldsymbol{\Sigma}_{(I+Q)\boldsymbol{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}}\boldsymbol{\Sigma^{-1}_{Z}}\boldsymbol{\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (68)
@@
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (69)
@@
Rank Deficiency Cases To avoid causing a rank deficiency of $I+Q$, we need to figure out the conditions that lead to rank deficiency. Here we show a simple case causing the rank deficiency. When the matrix $Q$ is a backward identity matrix [23],
@@
$\mathrm{HermiteForm}(I+Q)_{i}=\boldsymbol{0},\quad\exists i\in\left[1,k\right]$ (73)
where index $i$ is the row index, in this paper, the row index starts from $1$, and $\mathrm{HermiteForm}$ is the Hermite normal form [25].
Full Rank Cases Except for the rank deficiency cases, $I+Q$ has full rank and is invertible. Since $Q$ is a row equivalent to the identity matrix, we need to introduce the three types of elementary row operations as follows [52].
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case of adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise into the embeddings. To this end, we demonstrate that linear transform can be positive noise.

### v4

@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{split}&{|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\sigma^{2}_{Y_{k}}\end{bmatrix}-\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\bm{{\Sigma}_{Z+\epsilon}^{-1}}\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\mathrm{cov}^{2}(Y_{1},Z_{1}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{1}}}-\frac{\lambda}{\sigma^{4}_{Z_{1}}})&...&\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ \vdots&&\vdots\\ \mathrm{cov}(Y_{k},Z_{k}+\epsilon)\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&...&\sigma^{2}_{Y_{k}}-\mathrm{cov}^{2}(Y_{k},Z_{k}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{k}}}-\frac{\lambda}{\sigma^{4}_{Z_{k}}})\\ \end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&&\\ &\ddots&\\ &&\sigma^{2}_{Y_{k}}-\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}+\lambda\begin{bmatrix}\frac{1}{\sigma^{4}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{1},Z_{1})\mathrm{cov}(Y_{k},Z_{k})\\ \vdots&&\vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{k},Z_{k})\mathrm{cov}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{4}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}\right|\end{split}$
@@
We use the notation $\bm{v}=\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}^{T}$, and $\bm{V}=\mathrm{diag}(\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1}),\cdots,\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k}))$. And utilize the rule of determinants of sums Marcus (1990), then we have:
$\begin{split}\left|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}\Sigma^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}\right|=&\ \left|(\bm{\Sigma_{Y}-V})+\lambda\bm{v}\bm{v}^{T}\right|\\ =&\ \left|\bm{\Sigma_{Y}-V}\right|+\lambda\bm{v}^{T}(\bm{\Sigma_{Y}-V})^{-1}\bm{v}.\end{split}$ (42)
where $(\bm{\Sigma_{Y}-V})^{*}$ is the adjoint of the matrix $(\bm{\Sigma_{Y}-V})$. For simplicity, we can rewrite ${|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$ as:
@@
$\begin{split}&\left|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}\Sigma^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}\right|\\ =&\prod_{i=1}^{k}\left(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\cdot\frac{1}{\sigma^{2}_{Z_{i}}}\right)+\Omega\end{split}$ (43)
@@
where $\Omega=\lambda\cdot\bm{v}^{T}\cdot\bm{(\Sigma_{Y}-V)}^{*}\cdot\bm{v}$. The specific value of $\Omega$ can be obtained as:
@@
$\begin{split}\Omega=\lambda\cdot\begin{bmatrix}\dfrac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\dfrac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\cdot\begin{bmatrix}V_{11}&0&0\\ 0&\ddots&0\\ 0&0&V_{kk}\end{bmatrix}\cdot\begin{bmatrix}\dfrac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})\\ \vdots\\ \dfrac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\end{split}$ (44)
@@
$\begin{split}&{|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})\end{split}$ (47)
Accordingly, $|\bm{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|$ is:
$|\bm{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|=\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))$ (48)
As a result, $\frac{|\bm{\Sigma_{Y|Z+\epsilon}}|}{|\bm{\Sigma_{Y|Z}}|}$ is expressed as:
$\frac{|\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Y|Z+\epsilon}}|}=\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})}$ (49)
Combine Equations 49 and 39 together, the entropy change is expressed as:
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (50)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\bm{\epsilon})>0&\mathrm{if}\ M>0\\ \bigtriangleup S(\mathcal{T},\bm{\epsilon})\leq 0&\mathrm{if}\ M\leq 0\end{matrix}\right.$ (53)
@@ B.2 Add Gaussian Noise to Raw Images
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&H(\mathcal{T};\bm{X})-H(\mathcal{T};\bm{X+\epsilon})\\ =&\frac{1}{2}\log|\bm{\Sigma_{Y|X}}|-\frac{1}{2}\log|\bm{\Sigma_{Y|X+\epsilon}}|\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y|X}}|}{|\bm{\Sigma_{Y|X+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\bm{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\end{split}$ (55)
@@
$\bigtriangleup S(\mathcal{T},\bm{\epsilon})=\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}$ (56)
Clearly, the introduction of Gaussian noise to each pixel in the original images has a detrimental impact on the task. Note that some studies have empirically shown that adding Gaussian noise to partial pixels of input images may be beneficial to the learning task Li (2022) Zhang et al. (2023).
@@ C.1 Inject Linear Transform Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\bm{Z})=&H(\mathcal{T};\bm{Z})-H(\mathcal{T};\bm{Z}+Q\bm{Z})\\ =&H(\bm{Y};\bm{Z})-H(\bm{Z})-(H(\bm{Y};\bm{Z}+Q\bm{Z})-H(\bm{Z}))\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}+Q\bm{Z})\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma}_{\bm{Z}}||\bm{\Sigma_{Y}}-\bm{\Sigma_{YZ}}{\bm{\Sigma_{Z}^{-1}}}\bm{\Sigma_{ZY}}|}{|\bm{\Sigma}_{(I+Q)\bm{Z}}||\bm{\Sigma_{Y}-\Sigma_{YZ}}\bm{\Sigma^{-1}_{Z}}\bm{\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (57)
@@
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\bm{Z})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (58)
@@
Rank Deficiency Cases. To avoid causing a rank deficiency of $I+Q$, we need to figure out the conditions that lead to rank deficiency. Here we show a simple case causing the rank deficiency. When the matrix $Q$ is a backward identity matrix Horn & R. (2012),
@@
$\mathrm{HermiteForm}(I+Q)_{i}=\bm{0},\quad\exists i\in\left[1,k\right]$ (62)
where index $i$ is the row index, in this paper, the row index starts from $1$, and $\mathrm{HermiteForm}$ is the Hermite normal form Kannan & Bachem (1979).
Full Rank Cases. Except for the rank deficiency cases, $I+Q$ has full rank and is invertible. Since $Q$ is a row equivalent to the identity matrix, we need to introduce the three types of elementary row operations as follows Shores (2007).
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\bm{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case of adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise into the embeddings. To this end, we demonstrate that linear transform can be positive noise.